OpenStudy (kittiwitti1):

Anyone know the standard physical state of carbonate (CO3)?

OpenStudy (cuanchi):

aqueous? CO3^2-(aq)

OpenStudy (kittiwitti1):

@cuanchi I don't know. The total compound is this:\[SrCO_{3}\]

OpenStudy (kittiwitti1):

I know Strontium is in solid form and Carbon is solid, but Oxygen is a gas. Isn't carbonate a solid normally though?

OpenStudy (cuanchi):

in what topic are you talking about? as a compound in the nature? Strontium carbonate is a white, odorless, tasteless powder

OpenStudy (kittiwitti1):

Naturally occurring state, I suppose.

OpenStudy (luigi0210):

Also by solubility rules, most carbonates (unless with an alkali-metal) are insoluble

OpenStudy (kittiwitti1):

Er... maybe I'm not comprehending the problem. I'll get back to you guys on this...

OpenStudy (coolor):

you gotta specify the temperature and conditions tho

OpenStudy (coolor):

strontium carbonate is solid at room temp

OpenStudy (kittiwitti1):

@coolor @Luigi0210 @cuanchi it is this: http://prntscr.com/cwjcb3

OpenStudy (vincent-lyon.fr):

The "standard state of a substance at temperature T" is the most stable state of that substance under pressure \(P=10^5\) Pa, at temperature T. Water standard state at 300 K is liquid Water standard state at 400 K is vapour Water standard state at 250 K is ice-Ih https://en.wikipedia.org/wiki/Ice_Ih Carbon standard state at 300K is graphite (not diamond) Iron standard state at 300 K is solid alpha-iron (CCC) Strontium carbonate standard state at 300 K is solid (rhombic structure)

OpenStudy (vincent-lyon.fr):

If your strontium carbonate is in solution, the standard state is slightly more complicated to define. From English Wikipedia: _______________ For a substance in solution (solute), the standard state is the hypothetical state it would have at the standard state molality or amount concentration but exhibiting infinite-dilution behavior. The reason for this unusual definition is that the behavior of a solute at the limit of infinite dilution is described by equations which are very similar to the equations for ideal gases. Hence taking infinite-dilution behavior to be the standard state allows corrections for non-ideality to be made consistently for all the different solutes. Standard state molality is 1 mol kg−1, while standard state amount concentration is 1 mol dm−3. _________________

OpenStudy (caozeyuan):

your equation is incorrect is becuase there is no CO3(l), it should be \[Sr ^{2+}\left( aq \right)+CO _{3}^{2-}\left( aq \right)=SrCO _{3}\left( s \right)\]

OpenStudy (caozeyuan):

this euqation tells you that if you have Sr ion and carbonate ion in solution, you will get SrCO3 solid

OpenStudy (vincent-lyon.fr):

caozeyuan is correct: strontium carbonate being almost insoluble, the second part of my answer does not hold in this case. It would be valid, though, for solutes such as NaCl for instance.

OpenStudy (kittiwitti1):

the answer was to split up the \(CO_{3}\) guys, \((aq)\) is still wrong :\

OpenStudy (kittiwitti1):

*unbalanced equation:\[Sr(s)+C(s)+O_{2}(g)\rightarrow SrCO_{3}\]